what is the common difference in the numbers 1.05, 1.1, 1.15, 1.2?
Answer by swincher4391(1107)

What can you add to each term to get the next? 1.2-1.15 = .05, 1.15 - 1.1 = .05, 1.1-1.05 = .05.
.05 is your common difference.